# Dual Prototyping with Domain and Class Prototypes for Affective Brain-Computer Interface in Unseen Target Conditions

### v1

@@ VI Conclusion
This study proposes a novel pairwise learning framework with domain and class prototypes (PL-DCP) for EEG-based emotion recognition in unseen target conditions. Unlike existing transfer learning methods that require both source and target data for feature alignment, PL-DCP relies solely on source data for model training. Experimental results show that the proposed method achieves promising results even without using target domain data for training, with performance approaching or even surpassing some deep transfer learning models that heavily rely on target domain data. This suggests that combining feature disentanglement with domain and class prototypes helps generalize more reliable and stable characteristics of individual subjects. Additionally, the introduction of pairwise learning enhances the model’s resilience to label noise. These findings underscore the potential of this method for practical applications in aBCIs.
VII Acknowledgments
@@
This work was supported in part by the National Natural Science Foundation of China under Grant 62176089, 62276169 and 62201356, in part by the Natural Science Foundation of Hunan Province under Grant 2023JJ20024, in part by the Key Project of Xiangjiang Laboratory under Granted 23XJ02006, in part by the STI 2030-Major Projects 2021ZD0200500, in part by the Medical-Engineering Interdisciplinary Research Foundation of Shenzhen University under Grant 2024YG008, in part by the Shenzhen University-Lingnan University Joint Research Programme, and in part by Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions (2023SHIBS0003).

### v2

@@ 5.4 Conclusion
This study proposes a novel pairwise learning framework with domain and class prototypes (PL-DCP) for EEG-based emotion recognition in unseen target conditions. Unlike existing transfer learning methods that require both source and target data for feature alignment, PL-DCP relies solely on source data for model training. Experimental results show that the proposed method achieves promising results even without using target domain data for training, with performance approaching or even surpassing some deep transfer learning models that heavily rely on target domain data. This suggests that combining feature disentanglement with domain and class prototypes helps generalize more reliable and stable characteristics of individual subjects. Additionally, the introduction of pairwise learning enhances the anti-interference ability of the model to label noise. These findings underscore the potential of this method for practical applications inAffective Brain-Computer Interfaces(aBCIs). In the future work, we will further explore aBCIs models with better performance and more adaptability.
Acknowledgements
@@
This work was supported in part by the National Natural Science Foundation of China under Grant 62176089, 62276169 and 62201356, in part by the Natural Science Foundation of Hunan Province under Grant 2023JJ20024, in part by the Key Research and Development Project of Hunan Province under Grant 2025QK3008, in part by the Key Project of Xiangjiang Laboratory under Granted 23XJ02006, in part by the STI 2030-Major Projects 2021ZD0200500, in part by the Medical-Engineering Interdisciplinary Research Foundation of Shenzhen University under Grant 2024YG008, in part by the Shenzhen University-Lingnan University Joint Research Programme, and in part by Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions (2023SHIBS0003).